OpenStudy (anonymous):

simplify (5+x^-1)/(25-x^-2)

OpenStudy (anonymous):

first get rid of the stupid negative exponents so you can make sense out of this \[\frac{5+\frac{1}{x}}{25-\frac{1}{x^2}}\]

OpenStudy (anonymous):

then clear the fractions in one step by multiplying top and bottom by \[x^2\]

OpenStudy (anonymous):

if you want to be fancy you can factor the denominator as the difference of two squares and then cancel, but you will still have a compound fraction.

OpenStudy (anonymous):

\[\frac{5x^2+x}{25x^2-1}\] \[\frac{x(5x+1)}{(5x+1)(5x-1)}\] \[\frac{5}{5x+1}\]